The following data give information on the ages (in years) and the number of breakdowns during the past month for a sample ofWhat is the coefficient of determination R? for the sample? Round your answer to two decimal places. r2 =
Answer #1

Lease square regression equation.

\hat{y} = b0 + b1 *x

b0 is Intercept , we can find intercept using excel function =INTERCEPT( Y data set column range, x data set column range )

b1 is Slope , we can find slope using excel function =SLOPE( Y data set column range, x data set column range )

Age Breakdowns 21 =INTERCEPT(B2:38,A2:A8) |-1.18 =SLOPE(B2:38,A2:A8) 0.48

\hat{y} = -1.18 + 0.48 x

Similarly we can find correlation coefficient r using the function =CORREL( y, x )

11 Age Breakdowns minino =CORREL(B2:38,A2:A8) 10.97 7 15 8 19

r = 0.97

Therefore coefficient of determination r2 = 0.97*0.97

r2 = 0.94